40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 112 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  11537   Fri Aug 28 17:04:12 2015 SteveUpdatePEMGur interface box cable
Ara! Onara suru tsu-mori datta keh-do, un-chi ga de-chatta...
Quote:

 

Quote:

As reported previously, the transfer functions of the channels look fine. (i.e. All channels almost identical.)
I checked the chain from the unit input to the DAQ BNC connectors. They were all OK.

Today I have been checking the signals on the unit with the long DB37 cables connected.
I could not see anything on the Gur2 channels on the board. I looked at the DB37 for Gur2 and felt something is wrong.

I opened the housing of the cable and realized that all the pins are not fully inserted.
The wires were crimped improperly and prevents them to be fully inserted.

=> We need to redo the crimping to insert them.
=> We need to check the other side too.

I'm making a new long cable. Both connector ends of this X arm long cable were terrible. It was removed from the cable tray yesterday.

The Y arm Guralp is running fine. The interface box is open in front of 1X1 on a cart and it is alive!    Please be aware of it !

The new cable was made this way:

Pins were located with ohm meter for locations and both ends were cut off.The Belden 1424A cable than was soldered to DB37 and Guralp circular connector " IKPT06F16-26S-ND "

This cable will connect the ADC interface box CCD# DO 60506 to Guralp seismometer at the sout end.

The Guralp's each 3 axies will be connected through a twisted pair to the differential input op amp

Gur  ouput  vertical axis + -  on circular connector   A, B  to DB37 pin 1 &20

Gur output        N/S axis + -  on circular connector   C, D to DB37 pin 2 & 21

Gur output       E/W axis + -  on circular connector   E, F to DB37 pin 3 & 22

Power +12 VDC  from DB37 pin 29 to Gur circular connector pin c (lower case)

Power -12 VDC  from DB37 pin 24 to Gur circular connector pin  M Note: this connection was absent at the first test of this cable!

Ground              from DB37 pin 10 to Gur circular connector pin b (lower case)

Summery: I may destored the opamp  at the Guralp A at the south end.

 

 

Attachment 1: fixedNewLongCable.png
fixedNewLongCable.png
  11536   Fri Aug 28 02:20:35 2015 IgnacioUpdateLSCPRFPMI and MCL FF

A day late but here it is.

Eric and I turned on my SISO MCL Wiener filter elog:11535 during his PRFPMI 40min lock. We looked at the CARM_IN and CARM_OUT signals during the lock and with the MCL FF on/off. Here is the spectra:

  11535   Fri Aug 28 00:59:55 2015 IgnacioUpdateIOOFinal SISO FF Wiener Filter for MCL

This is my final SISO Wiener filter for MCL that uses the T240-X seismo as its witness.

The main difference between this filter and the one on elog:11532 is the actual 1/f rolloff this filter pocesses. My last filter had a pair of complex zeroes at 2kHz, that gave the filter some unusual behavior at high frequencies, thanks Vectfit. This filter has 10 poles and 8 zeroes, something Vectfit doesn't allow for and needs to be done manually.

The nice thing about this filter is the fact that Eric and I turned this filter on during his 40 min PRFPMI lock last night, Spectra for this is coming soon.

This filter lives on the static Wiener path on the OAF machine, MCL to MC2, filter bank 7.

Anyways, the usual plots are shown below. 

 

Filter:

T240-X (SISO)

 

Training data + Predicted FIR and IIR subtraction:

 

Online subtraction results:(High freq. stuff shown for noise injection evaluation of the filter)

MCL

YARM

Subtraction performace:

  11534   Thu Aug 27 04:23:04 2015 ericqUpdateLSCPRFPMI is back

Got to a 40 minute lock tonight. All other locks broke because of me poking something. 

I redid some sensing excitations, right after carefully measuring the CARM OLG at its excitation frequency, so I can get at the open loop PD response. 

I also used a MCL feedforward filter of Ignacio's which did not inject any observable noise into the CARM error signal during PRFPMI lock. He will make some elog about this. 

  11533   Thu Aug 27 02:09:14 2015 ericqUpdateLSCAUX X Laser Current Changed

I spent some time tonight chasing down the cause of huge RIN in the X green PDH transmitted light, which I had started seeing on Monday. This was preventing robust locking, since the ALS sensing noise was ~10x worse above 50Hz, thus making the AO transition much flakier (though, impressively, not impossible!)

I went down to the X end, and found that turning the laser diode current down by 0.1A (from 2.0 to 1.9) smoothed things out completely. Unfortunately, this causes the power to drop, from GTRX of 0.45 to 0.3, but the ALSX sensitivity is unchanged, as compared with the recenent "out of loop" template. 

This also seems to have changed the temperatures of the good modes, as no beat was evident at the previously good temperature. Beats were found at +5400 and +10500 counts on the slow servo offset slider; I suspect the third lies around the edge of the DAC range which is why I couldn't uncover it. In any case, I've parked it at 10500 for now, and will continue locking; nailing it down more precisely and offloading the slider offset to the laser controller will happen during daytime work...

  11532   Thu Aug 27 01:41:41 2015 IgnacioUpdateIOOTriply Improved SISO (T240-X) FF of MCL

Earlier today I constructed yet another SISO filter for MCL. The one thing that stands out about this filter is its strong roll off wink. This prevents high frequency noise injection into YARM. The caviat, filter performance suffered broken heart quite a bit, but there is subtraction going on.

I have realized that Vectfit lacks the ability of constraining the fits it produces, (AC coupling, rolloff, etc) even with very nitpicky weighting. So the way I used vectfit to produce this filter will be explained in a future eLOG, I think it might be promising. 

Anyways, the usual plots are shown below. 

 

Filter:

T240-X (SISO)

 

 

Training data + Predicted FIR and IIR subtraction:

 

Online subtraction results:(High freq. stuff shown for noise injection evaluation of the filter)

MCL
 
YARM
 
 
 

Subtraction performace:

 

  11531   Tue Aug 25 16:39:06 2015 SteveUpdatesafetysafety training

Alessandra Marrocchesi received 40m specific basic safety training yesteday.

  11530   Tue Aug 25 16:33:31 2015 Ignacio, SteveConfigurationPEMSeismometer enclosure copper foil progress

Steve ordered about two weeks ago a roll of 0.5 mm thick copper foil to be used for the inside of the seismometer cans. The foil was then waterjet cut by someone in Burbank to the right dimensions (in two pieces, a side and a bottom for each of the three cans).

Today, we glued the copper foil (sides only) inside the three seismometer cans. We used HYSOL EE4215/HD3561(Data Sheet) as our glue. It is a "high impact, low viscocity, room temperature cure casting" that offers "improved thermal conductivity and increased resistance to heat and thermal shock." According to Steve, this is used in electronic boards to glue components when you want it to be thermal conductive.

We are going to finish this off tomorrow by gluing the bottom foil to the cans. The step after this involves soldering the side to the bottow and where the side connects. We have realized that the thermal conductivity of the solder that we are using is only ~50. This is 8 times smaller than that of copper and wil probably limit how good a temperature gradient we will have.

Some action shots,

  11529   Tue Aug 25 16:09:54 2015 ericqUpdateIOOIMC Tweak

A little more information about the IMC loop tweak...

I increased the overall IMC loop gain by 4dB, and decreased the FAST gain (which determines the PZT/EOM crossover) by 3dB. This changed the AO transfer function from the blue trace to the green trace in the first plot. This changed the CARM loop open loop TF shape from the unfortunate blue shape to the more pleasing green shape in the second plot. The red trace is the addition of one super boost. 

 

Oddly, these transfer functions look a bit different than what I measured in March (ELOG 11167), which itself differed from the shaping done December of 2014 (ELOG 10841). 


I haven't yet attempted any 1F handoff of the PRMI since relocking, but back when Jenne and I did so in April, the lock was definitely less stable. My suspicion is that we may need more CARM supression; we never computed the loop gain requirement that ensures that the residual CARM fluctuations witnessed by, say, REFL55 are small enough to use as a reliable PRMI sensor.

I should be able to come up with this with data from last night. 

Attachment 1: imcTweak.pdf
imcTweak.pdf
Attachment 2: CARM_TF.pdf
CARM_TF.pdf
  11528   Tue Aug 25 04:15:51 2015 ericqUpdateLSCPRFPMI is back

More PRFPMI locks tonight. Right now, it's been locked for 22+ minutes, though with the PRMI still on 3F signals. I think the MC2/AO crossover needs some reshaping; there's a whole bunch of noise injected into CARM around 600 Hz, which is where the two paths differ by 180deg. (Addendum: broke lock at ~27 minutes, 4:16AM)

For most of this lock, sensing matrix excitations have been running for daytime analysis. 

The nominal IMC loop gain / EOM crossover were making the AO path very marginal. I've adjusted the nominal settings and autolocker scripts. 

There was some weird behavior of X green PDH earlier... Broadband RIN seen in ALS-TRX, coherent with the DC output of the beat PD, so really on the light. I fiddled with the end setup, and it mostly went away, though I didn't intentionally change anything. Disconcerting. 

  11527   Mon Aug 24 16:46:49 2015 ericqUpdateSUSETMY Oplev laser power is falling

Repeated for all optics, ETMY seems like the only one sharply dropping for now (PRM is all over the place and hard to gauge, since we often leave it partially- or mis-aligned):

 


Hacky bits:

Bash:

find /opt/rtcds/caltech/c1/burt/autoburt/snapshots/2015 -wholename "*00:0*"
| xargs ack --nogroup "OL_SUM_OUT16 1"
| grep -v 'SUS-MC'
| sed -e 's/.*2015/2015/g' -e 's/\/c1.*C1:SUS-/, /g' -e 's/_OL.*\([0-9]\..*$\)/, \1/g' -e 's/\//-/g'
| sort | uniq > allOL.txt

qontrols@pianosa|~ > head allOL.txt 
2015-Apr-10-00:07, BS, 1.146766113281250e+03
2015-Apr-10-00:07, ETMX, 1.597261328125000e+04
2015-Apr-10-00:07, ETMY, 4.331762207031250e+03
2015-Apr-10-00:07, ITMX, 6.488521484375000e+03
2015-Apr-10-00:07, ITMY, 1.387590234375000e+04
2015-Apr-10-00:07, PRM, 8.352053833007812e+02
2015-Apr-10-00:07, SRM, 6.099560928344727e+01
2015-Apr-1-00:07, BS, 1.180478149414062e+03
2015-Apr-1-00:07, ETMX, 1.584842480468750e+04

Python:

 

olsum = pd.read_csv('allOL.txt',parse_dates=True, names=['Date','Optic','Sum'])
olsum['Date'] = pd.to_datetime(olsum['Date']) # Automatic parsing didn't work for some reason
olpivot = olsum.pivot(index='Date',columns='Optic',values='Sum')
olpivot.plot()
Attachment 1: OLsum_trend.pdf
OLsum_trend.pdf
  11526   Mon Aug 24 16:10:07 2015 ericqUpdateSUSETMY Oplev laser power is falling

Today I noticed the box around the ETMY oplev sum flashing red, as it dipped below 1k. I don't recall seeing this recently, so I wanted to look up the history.

However, we've been having trouble with our minute (and longer) trend data, so I had to hack it out a bit... Here is the unfortunate result:

I think we can be fairly confident that this is not due to alignment drifts, we generally keep the QPD reasonably well centered. I also recentered it today, and the counts remained at ~1k. 


Details of the hack that got me this data:

I ended up looking at the BURT snapshots from every night at midnight, which report a number for ETMY_OL_SUM_OUT16, and making a text file with dates and values with the following BASH spaghetti:

find /opt/rtcds/caltech/c1/burt/autoburt/snapshots/2015 -wholename "*00*/*scy*" |xargs ack --nogroup "ETMY_OL_SUM_OUT16 1" |  sed -e 's/.*2015/2015/g' -e 's/\/c1.*\([0-9]\..*$\)/, \1/g' -e 's/\//-/g'  > ETMYsum.txt

This produces a file full of unsorted lines like: 2015-Aug-23-00:07, 1.106459228515625e+03

The python package pandas is good at parsing dates and automatically plotting time series: 

olsum = pandas.read_csv('ETMYsum.txt', index_col=0, parse_dates=True)
olsum.plot()
Attachment 1: ETMYsum_trend.pdf
ETMYsum_trend.pdf
  11525   Mon Aug 24 14:05:47 2015 ericqUpdateSUSEricG Investigating L2A

This afternoon, I showed Eric Gustafson some of the basics of making swept sine measurements with DTT. We turned off the f2a filters and oplev damping on the BS and made a cursory measurement of the transfer function from position drive to the oplev signals. 

He will be in the lab periodically to continue this line of investigations. 

  11524   Sat Aug 22 15:48:32 2015 KojiSummaryLSCArm locking recovery

As per Ignacio's request, I restored the arm locking.

- MC WFS relief

- Slow DC restored to ~0V

- Turned off DARM/CARM

- XARM/YARM turned on

- XARM/YARM ASS& Offset offloading

  11523   Fri Aug 21 17:15:13 2015 SteveUpdatePEMGur interface box

 

Quote:

As reported previously, the transfer functions of the channels look fine. (i.e. All channels almost identical.)
I checked the chain from the unit input to the DAQ BNC connectors. They were all OK.

Today I have been checking the signals on the unit with the long DB37 cables connected.
I could not see anything on the Gur2 channels on the board. I looked at the DB37 for Gur2 and felt something is wrong.

I opened the housing of the cable and realized that all the pins are not fully inserted.
The wires were crimped improperly and prevents them to be fully inserted.

=> We need to redo the crimping to insert them.
=> We need to check the other side too.

I'm making a new long cable. Both connector ends of this X arm long cable were terrible. It was removed from the cable tray yesterday.

The Y arm Guralp is running fine. The interface box is open in front of 1X1 on a cart and it is alive!    Please be aware of it !

  11522   Fri Aug 21 08:38:37 2015 SteveUpdateIOOPMC locked

PMC needed to be locked manually.cheeky

Attachment 1: PMClocked.png
PMClocked.png
  11521   Thu Aug 20 18:08:28 2015 IgnacioFrogs40m upgradingFatality. Something broke.

So I made coffee at 1547 and was astonished to find the above. Its a sad, very sad day.

At first I thought that something (a gravity wave?) or someone, accidentally hit the thing and it fell and broke. But Koji told me that the janitor was cleaning around the thing and it did indeed fell accidentally.

  11520   Thu Aug 20 11:31:55 2015 KojiUpdatePEMGur interface box

As reported previously, the transfer functions of the channels look fine. (i.e. All channels almost identical.)
I checked the chain from the unit input to the DAQ BNC connectors. They were all OK.

Today I have been checking the signals on the unit with the long DB37 cables connected.
I could not see anything on the Gur2 channels on the board. I looked at the DB37 for Gur2 and felt something is wrong.

I opened the housing of the cable and realized that all the pins are not fully inserted.
The wires were crimped improperly and prevents them to be fully inserted.

=> We need to redo the crimping to insert them.
=> We need to check the other side too.

Attachment 1: IMG_1958.JPG
IMG_1958.JPG
  11519   Thu Aug 20 11:09:10 2015 ranaUpdateIOOsome points about seismic FF
  • When plotting the subtraction performance, we mainly care about the 0.5 - 10 Hz band, so we care about the RMS in this band. Don't integrate over the whole band.
  • When calculating the Wiener filter, you must use the pre-weighting so as to not let the Wiener residual be dominated by the out of band signals. We don't want the filter to try to do anything outside of the 0.5 - 10 Hz band.
  • Somehow, we want to assign a penalty for the filter to have high frequency gain. We do NOT want to slap on an ad-hoc low pass filter. The point of the Wiener filtering is to make the optimum.
  • What is the reason for the poor filter performance from 0.5 - 2 Hz ? If we use the frequency domain (Dmass) subtraction technique, we can do better, so there's some inefficiency in this process.
  • we're getting too much of the 3 Hz stack mode coupling into MCL. I think this means that our damping filters should be using RG around the suspension eigenmodes rather than just simple velocity damping. We had this years ago, but it caused some weird interaction with the angular loops...to be puzzled out.
  11518   Thu Aug 20 02:31:09 2015 ericqUpdateLSCPRFPMI is back

PRFPMI locking has been revived.

I've had 6 5min+ locks so far; arm powers usually hit ~125 for a recycling gain of about 7; visibility is about 75%

The locking script takes a little under 4 minutes to take you from POX/POY lock to PRFPMI if you don't have to stop and adjust anything.

At Koji's suggestion, I used digital REFL11 instead of CM_SLOW, which got me to a semistable lock with some RF, at which time I could check the CM_SLOW situtation. It seemed like the whitening Binary IO switch got out of sync with the digital FM status somehow... 

I've been making the neccesary changes to the carm_cm_up script. I also added a small script which uses the magnitude of the I and Q signals to set the phase tracker gain automatically based on some algebra Koji posted in an ELOG some years ago. 

The RF transition seems much smoother now, most likely due to the improved PRC and ALS stability. In fact, it is possible to hold at arm powers of >100 solely on the digital servos; I don't think we were able to do this before until the AO had kicked in. 

Right now I'm losing lock when trying to engage the CARM super boost. I also haven't switched the PRMI over to 1F signals yet. Would be good to hook the SR785 back up for a loop TF, but I'll stop here for tonight since our SURFs are presenting bright and early tomorrow morning. 

Attachment 1: lock.pdf
lock.pdf
  11517   Wed Aug 19 07:58:25 2015 SteveUpdatePEMGur interface box

Koji and Steve,

We took transferfunctions of each channel yesterday. They were identical ?. I will check the cables from ADC to DAQ next.

Attachment 1: GurADCbox.jpg
GurADCbox.jpg
  11516   Wed Aug 19 01:45:10 2015 IgnacioUpdateIOODoubly Improved SISO (T240-X) FF of MCL

Today I tried and doubly-improved SISO FF filter on MCL. This filter has a stronger rolloff than the previous SISO filters I have tried. The rolloff most definelty helped towards reducing the ammount of noise being injected into YARM. Below is the usual stuff:

 

Filter:

T240-X (SISO)

 

 

Training data + Predicted FIR and IIR subtraction:

 

Online subtraction results:

MCL
YARM
MCL TRANS
 
 
 

Subtraction performace:

  11515   Wed Aug 19 00:55:35 2015 IgnacioUpdateLSCLSC-YARM-EXC to LSC-YARM-IN1 TF measurement + error analysis

Yesterday, Rana, Jessica and I measured the Transfer function from LSC-YARM-EXC to LSC-YARM-IN1. 

The plot below shows the magnitude and the phase of the measured transfer function. It also shows the normalized standard error in the estimated transfer function magnitude; the same quantity can be applied to the phase, only in this case it is interpreted as its standard deviation (not normalized). It is given by

 \frac{[1-\gamma_{xy}^2(f)]^{1/2}}{|\gamma_{xy}(f)|\sqrt{2n_{d}}}

where \gamma_{xy}^2(f) is the ordinary coherence function and n_{d} is the number of averages used at each point of the estimate, in the case here we used 9 averages. This quantity is of interest to us in order to understand how the accuracy of transfer function measurement affects the ammount of subtraction that can be achieved online.

 

Since this transfer function is flat from 1-10 Hz (out of phase by 180 deg), this means that we can apply our IIR wiener filters direclty into YARM without taking into account the TF by prefiltering our witnesses with it. Of course this is not the case if we care about subtractions at frequencies higher than 10 Hz, but since we are dealing with seismic noise this is not a concern.

The coherence for this transfer function measurement is shown below,

  11514   Tue Aug 18 11:16:17 2015 SteveUpdatePEMGur interface box is wonky

The Guralp ADC  interface box  D060506 is ready for inspection. It is in front of 1X1 with open top and running.cool

Obviously c7 as 1 miroF cap is missing.blush

Quote:

Let's dismantle the I/F unit from the rack and connect the cable with the lid open.
We need to trace the signal.

 

Attachment 1: IMG_0009.JPG
IMG_0009.JPG
  11513   Tue Aug 18 03:56:09 2015 ericqUpdateLSClocking efforts

Now that the updated ALS is stable, and the PRC angular FF is revived, I've been working on relocking PRFPMI. While the RMS arm fluctuations are surely smaller than they used to be, there is no noticible difference to the ears when buzzing around resonance, but this doesn't really mean much. 

Frustratingly, I am not able to stably blend in any RF CARM error signal into the slow length control path (i.e. CARM_B). Bringing AS55 Q into DARM with the 20:0 integrator is working fine, but we really need to supress CARM to get anywhere. I'm not sure why this isn't working; poking around into the settings that were used when we were regularly locking didn't turn up any differences as far as I could tell. Investigations continue... 

Some minor changes to the locking script were made, to account for the increased ALS displacement sensitivity from the longer delay line. 


Since the ALS is now in a fairly stable state, I've updated the calibrated PSD template at /users/Templates/ALS/ALS_outOfLoop_Ref.xml, and added some coherence plots for some commonly coupled quantities (beat signal amplitude, IR error signal, green PDH error signal and green transmission). 

Attachment 1: newALSref.pdf
newALSref.pdf
Attachment 2: xCoh.pdf
xCoh.pdf
Attachment 3: yCoh.pdf
yCoh.pdf
  11512   Mon Aug 17 17:48:12 2015 KojiUpdatePEMWasps obliterated maybe...

We found the same wasp in the 40m. Megan found it walking behind Steve desk!

  11511   Sun Aug 16 23:26:40 2015 EveUpdateGeneralGaussianity tests

I've continued to work on my Gaussianity tests for S5 L1 data. 

Following the statistical measure in Ando et al. (2003), I've calculated the Laguerre coefficient, c2, for all frequencies present in my S5 L1 PSD as a metric of Gaussianity. When c2 is zero, the distribution is Gaussian. A positive c2 corresponds to glitchy noise, while a negative c2 suggests stationary noise.

Below is a plot displaying variation in c2 for this PSD:

By observing the c2 value and histogram of distribution of various PSD values at a given frequency, we can elucidate statistical differences in the Gaussian nature of noise at that frequency which are unclear in the standard PSD.

Attachment 1: Gaussianity_noc1.png
Gaussianity_noc1.png
  11510   Sat Aug 15 02:10:35 2015 IgnacioUpdateLSCMCL FF => YARM FF

In my last post I calculated the different subtractions (offline) that could be done to YARM alone just to get a sense of what seismometers were better witnesses for the Wiener filter calculation. 

In this eLOG I show what subtractions can be done when the MCL has FF on (as well as Eric's PRC FF), with the SISO filter described on elog:11496.

The plot below shows what can be done offline,

What is great about this results is that the T240-X and T240-Y channels are plenty enough to mitigate any remaining YARM seismic noise but also to get rid of that nasty peak at 55 Hz induced by the MCL FF filter.

The caviat, I haven't measured the TF for the ETMY actuator to YARM control signal. I need to do this and recompute the FIR filters with the prefiltered witnesses in order to move on to the IIR converions and online FF!

 

Attachment 1: YARM_LIVES.png
YARM_LIVES.png
  11509   Fri Aug 14 23:49:34 2015 KojiSummaryGeneralB&K Shaker fixed

I fixed a shaker that was claimed to be broken. I had to cut the rubber membrane to open the head.

Once it was opened, the cause of the trouble was obvious. The soldering joint could not put up with the motion of the head.

It is interesting to see that the spring has the damping layer between the metal sheets.

After the repair the DC resistance was measured. It was 1.9Ohm. The side of the shaker chassis said "3.5Ohm, Max 15VA". So it can take more than 4A (wow).

I gave 2A DC from the bench top supply and turn the current on and off. I could confirm the head was moving.

I'll claim the use of this shaker for the seismometer development.

Attachment 1: IMG_1947.JPG
IMG_1947.JPG
  11508   Fri Aug 14 21:40:26 2015 IgnacioUpdateLSCQuick static offline subtractions of YARM

Plotte below are the resultant subtractions for YARM using different witness configurations,

The best subtraction happens with all the channels of both the GUR1 and T240 seismometers, but one gets just as good subtraction without using the z channels as witnesses. 

Also, why is the T240 seismometer better at subtracting noise for YARM compared to what GUR1 alone can acomplish? Using only the X and Y channels for the T240 gave the third best subtraction(purple trace). 

My plan for now is as follows:

1) Measure the transfer function from the ETMY actuator to the YARM control signal

2) Collect data for YARM when FF for MCL is on in order to see what kind of subtractions can be done.

Attachment 1: arms_wiener.png
arms_wiener.png
  11507   Fri Aug 14 17:20:01 2015 JenneUpdatePEMGur interface box is wonky

IIRC, the Guralp box's 3rd set of channels do not have all of the modifications that were made on channels 1 and 2.

  11506   Fri Aug 14 12:10:08 2015 KojiUpdatePEMGur interface box is wonky

Let's dismantle the I/F unit from the rack and connect the cable with the lid open.
We need to trace the signal.
 

  11505   Fri Aug 14 09:07:45 2015 SteveUpdatePEMGur interface box is wonky

Atm1, Before cable swap

Atm2, The long cables were swapped at the input of the interface box.

We can conclude that the problem is in the interface box

I wonder if interface box input 3 is wired?

Attachment 1: Gur1&2.png
Gur1&2.png
Attachment 2: longCablesSwapped.png
longCablesSwapped.png
  11504   Thu Aug 13 23:57:33 2015 IgnacioUpdateLSCYARM coherence plots

I took data from 1123495750 to 1123498750 GPS time (Aug 13 at 3AM, 50 mins of data) for  C1:LSC-YARM_OUT_DQ, and all T240 and GUR1 channels.

Here is the PSD of the YARM_OUT, showing the data that I will use to train the FIR filter:

Coherence plots for YARM and all channels of T240 and GUR1 sesimometers are shown below. This will help determine what regions to preweight the best before computing FIR filter. They also show how GUR1 is back to work compared to those of elog:11457.

 

 

Attachment 1: YARM_psd.png
YARM_psd.png
Attachment 2: YARM_GUR1_COH.png
YARM_GUR1_COH.png
Attachment 3: YARM_STS_COH.png
YARM_STS_COH.png
Attachment 4: YARM_GUR1_COH.png
YARM_GUR1_COH.png
  11503   Thu Aug 13 20:32:07 2015 IgnacioUpdateLSCWorking towards YARM FF

The mode cleaner FF static filtering is by no means done. More work has to be done in order to succefuly implement it, by the means of fine tuning the IIR fit and finding better MISO Wiener filters. 

I have begun to look at implementing FF to the YARM cavity for several reasons.

1) Even if the mode cleaner FF is set up as best as we can, there will still be seismic noise coupling into the arm cavities.

2) YARM is in the way of the beam path. When locking the IFO, one locks YARM first, then XARM. This means that it makes sense to look at YARM FF first rather than XARM.

3) XARM FF can't be done now since GUR2 is sketchy.

I'm planning on using this eLOG entry to document my Journey and Adventures (Chapter 2: YARM) to the OPTIMAL land of zero-seismic-noise (ZSN) at the 40m telescope.

 

  11502   Thu Aug 13 12:06:39 2015 Jessica SummaryIOOBetter predicted subtraction did not work as well Online

Yesterday I adjusted the preweighting of my IIR fit to the transfer function of MC2, and also managed to reduce the number of poles and zeros from 8 to 6, giving a smoother rolloff. The bode plots are pictured here:

The predicted IIR subtraction was very close to the predicted FIR subtraction, so I thought these coefficients would lead to a better online filter.

However, the actual subtraction of the MCL was not as good and noise was injected into the Y arm.

The final comparison of the subtraction factors between the online and offline data showed that the preweighting, while it improved the offline subtraction, needs more work to improve the online subtraction also.

Attachment 1: newBodeX.png
newBodeX.png
Attachment 2: newBodeY.png
newBodeY.png
Attachment 3: pred_Sub.png
pred_Sub.png
Attachment 4: MCLSub.png
MCLSub.png
Attachment 5: YarmSub.png
YarmSub.png
Attachment 6: comparison.png
comparison.png
  11501   Wed Aug 12 22:33:36 2015 IgnacioUpdateIOORe-measured MC2 -> MCL TF

Since I will need to do transfer function measurements in order to implement FF for the arms and the MC2's yaw and pitch channels, I decided to practice this by replicating the transfer function measurement Eric did for MC2 to MCL. I followed his procedure and the data that I aquired for the TF looked as shown below,

About five minutes of data were taken (0.05 Hz resolution, 25 averages) by injecting noise from 1 to 100 Hz. The TF coherence looked as below,

Attachment 1: bode_TF.png
bode_TF.png
Attachment 2: Coherence.png
Coherence.png
  11500   Wed Aug 12 16:48:26 2015 IgnacioUpdateIOOBetter? Nope. MISO WIener (T240-X and T240-Y) FF of MCL

Last night, I also worked on a "better" (an improvement, I thought) of the MISO Wiener filter (T240-X and T240-Y witnesses) IIR filter. The FF results are shown below:

Filter:

T240-X

T240-Y

 

Training data + Predicted FIR and IIR subtraction:

Online subtraction results:

MCL
YARM

Subtraction performace:

 Although the predicted FIR and IIR results are "better" than the previous MISO filter, the subtraction performance for this filter is marginally better if not worse (both peak at 15 dB, in slightly different regions). 

Attachment 1: stsx.png
stsx.png
Attachment 2: stsy.png
stsy.png
Attachment 3: performance.png
performance.png
Attachment 4: mcliir.png
mcliir.png
Attachment 5: yarmiir.png
yarmiir.png
Attachment 6: sub.png
sub.png
  11499   Wed Aug 12 16:39:46 2015 IgnacioUpdateIOOMISO WIener (T240-X and T240-Y) FF of MCL

Last night I performed some MISO FF on MCL using the T240-X and T240-Y as witnesses. Here are the results:

Filter:

T240-X

T240-Y

 

Training data + Predicted FIR and IIR subtraction:

Online subtraction results:

MCL
YARM

Subtraction performace:

Attachment 1: stsx.png
stsx.png
Attachment 2: stsy.png
stsy.png
Attachment 3: performance.png
performance.png
Attachment 4: sub.png
sub.png
Attachment 5: mcliir.png
mcliir.png
Attachment 6: yarmiir.png
yarmiir.png
  11498   Wed Aug 12 14:35:46 2015 ericqUpdateComputer Scripts / ProgramsPDFs in ELOG

I've tweaked the ELOG code to allow uploading of PDFs by drag-and-drop into the main editor window. Once again we can bask in the glory of 

(You may have to clear your browser's cache to load the new javascript)

Attachment 1: smooth.pdf
smooth.pdf
  11497   Wed Aug 12 11:44:13 2015 ericqUpdatePEMGur2 Channels still wonky

In previous elogs, we saw that the X and Y spectra out of GUR2 (X end Guralp seismometer) looked strange (i.e. inconsistent with the GUR1 spectra). 

This morning, Steve and I brought the handheld control unit to the Guralp to center the test mass, by adjusting the centering potentiometers inside the unit while monitoring the voltage readout corresponding to the DC mass position (manual has instructions). 

At first glance, this seemed like the likely culprit, as the offsets for the horizontal directions were much larger than the vertical one. We zeroed all three to within a mV or two. Unfortuntately, the spectra look the same as they did 10 hours ago. no

Since we already had the kit out, we checked the offsets for GUR1. Only the "East/West" had an offset over 50mV, so we zeroed that one, but left the others alone. 

Attachment 1: gur2_centered.png
gur2_centered.png
  11496   Wed Aug 12 01:32:18 2015 IgnacioUpdateIOOImproved SISO (T240-X) FF of MCL

In my previous elog:11492, I stated that in order to improve the subtraction and reduce the injection of high frequency noise we want the filter's magnitude to have a 1/f rolloff.

I implemented this scheme on the filter SISO filter previously analyzed. The results are shown below.

The filters bode plot:

The nice 1/f rollof is the main change here. Everything else remained pretty much the same.

The predicted FIR and IIR subtractions:

Everything looks right but that hump at 8 Hz. I used 8 pairs of poles/zeros to get this subtraction.

The online MCL subtraction:

This looks better than I expected. One has to keep in mind that I ran this at 1 AM. I wonder how well this filter will do during the noisier hours of the day. The RMS at high frequencies doesn't look great, there will definitely be noise being injected into the YARM signal at high frequencies.

Measuring the YARM signal:

There is still noise being injected on YARM but it is definitely much better than the previous filter. I'm thinking about doing some IIR subtraction on the arms now to see if I can get rid of the noise that is being injected that way, but before I embark on that quest I will rething my prefiltering.

The plot below shows the ratio of the unfiltered versus filtered ASDs for the FIR and IIR subtraction predictions as well as for the measured online IIR subtraction. Positive dB means better subtraction.

Attachment 1: filter.png
filter.png
Attachment 2: stsx.png
stsx.png
Attachment 3: mclonline.png
mclonline.png
Attachment 4: yarmonline.png
yarmonline.png
Attachment 5: sub.png
sub.png
  11495   Tue Aug 11 18:43:42 2015 JessicaUpdateIOOMCL Online Subtraction

Today I finished fitting the transfer function to a vectfit model for seismometers T240_X and T240_Y, and then used these to filter noise online from the mode cleaner. 

The Bode plot for T240_X is in figure 1, and T240_Y is in figure 2. I made sure to weight the edges of the fit so that no DC coupling or excessive injection of high frequency noise occurs at the edges of the fit.

I used C1:IOO-MC_L_DQ as the first channel I filtered, with C1:IOO-MC_L_DQ(RMS) for RMS data. I took reference data first, without my filter on. I then turned the filter on and took data from the same channel again. The filtered data, plotted in red, subtracted from the reference and did not inject noise anywhere in the mode cleaner. 

I also looked at C1:LSC-YARM_OUT_DQ and C1:LSC-YARM_OUT_DQ(RMS) for its RMS to see if noise was being injected into the Y-Arm when my filter was implemented. I took reference data here also, shown in blue, and compared it to data taken with the filter on. My filter, in pink, subtracted from the Y-Arm and injected no noise in the region up to 10 Hz, and only minimal noise at frequencies ~80 Hz. Frequencies this high are noisy and difficult to filter anyways, so the noise injection was minimal in the Y-Arm. 

Attachment 1: SeisX_bode.png
SeisX_bode.png
Attachment 2: SeisY_bode.png
SeisY_bode.png
Attachment 3: MCL_first.png
MCL_first.png
Attachment 4: Yarm_first.png
Yarm_first.png
  11494   Tue Aug 11 16:13:28 2015 EveUpdateGeneralGaussianity tests

I’m working on a code to determine the Gaussianity of a PSD.

 

Motivation:

It can be difficult to distinguish between GW events and non-Gaussian noise, especially in burst searches. By characterizing noise Gaussianity, we can better recognize noise patterns and distinguish between GW events and noise.

 

What I did:

I analyzed an hour of S5 L1 data. First, I plotted a timeseries, just to see what I was working with. Then, I produced a PSD (technically, an ASD) for the timeseries using Welch’s method in Python.

 

I split the data segment into smaller time-chunks and then produced a PSD for each chunk. All PSDs were superimposed in one plot. Here’s a plot for 201 time-chunks of equal length:

For a specific frequency, I can view the spread in PSD value through the production of a histogram.

 

Results:

I’ve made histograms displaying varying PSD values for the 201 PSD plot at 100 Hz, 500 Hz, and 1kHz.

For Gaussian noise, an exponential decay plot is expected. I will continue this analysis by following the statistical method in Ando et al. 2003 to calculate specific values indicative of the Gaussianity of various distributions. I’ll then look at different periods of time in the S5 L1 data to find periods of time suggesting non-Gaussian behavior.

  11493   Tue Aug 11 11:56:36 2015 Ignacio, JessicaUpdatePEMWasps obliterated maybe...

The wasp terminator came in today. He obliterated the known wasp nest.

We discovered a second wasp nest, right next to the previous one...

Jessica wasn't too happy the wasps weren't gone!

  11492   Tue Aug 11 11:30:19 2015 IgnacioUpdateIOOSISO (T240-X) FF of MCL

Last night we finally got some online subtraction going. The filter used is described in the post this eLOG is @eLOG 11488

The results were as follow:

The filter worked as expected when subtracting noise out of MCL,

There is about a factor of 6 subtraction at the ~3Hz resonant peak. The static IIR filter predicted a factor of 6-7 subtraction of this peak as well.

The 1.2 Hz resenonant feature improved by a factor of 3. This should improve quite drastically when I implement the y-channel of the T240 seismo.

There is some high frequency noise being injected, not very noticeable, but present. 

We then took a look at the power in the MC when the filter was on,

The power being transmitted in the cavity was not as stable as with the feedforward on. We believe that the filter is not at fault for this as Eric mentioned to me that the MC2 actuator lacked some sort of compensation that I need to understand a bit better.

YARM was then locked when the filter was on and we took a look at how it was doing. There was stationary sound arising from the locking of the YARM, leading us to believe that the filter might have injected some noise in the signal. IT DID.

The filter injected nasty high frequency noise at YARM from 11 Hz and on. This is to be expected since the filter did not roll off to zero at high frequencies. Implementing a 1/f rolloff should mitigate some of the injected noise.

 Also, as one can see above, subtraction by around a factor of 2 or so, was induced by the mode cleaner feedforward subtraction.

Attachment 1: MCL.png
MCL.png
Attachment 2: MCTRANS.png
MCTRANS.png
Attachment 3: YARM.png
YARM.png
  11491   Tue Aug 11 10:13:32 2015 JessicaUpdateGeneralConductive SMAs seem to work best

After testing both the Conductive and Isolated front panels on the ALS delay line box using the actual beatbox and comparing this to the previous setup, I found that the conductive SMAs improved crosstalk the most. Also, as the old cables were 30m and the new ones are 50m, Eric gave me a conversion factor to apply to the new cables to normalize the comparison.

I used an amplitude of 1.41 Vpp and drove the following frequencies through each cable:
      X: 30.019 MHz          Y: 30.019203 MHz

which gave a difference of 203 Hz. 

In the first figure, it can be seen that, for the old setup with the 30m cables, in both cables there is a spike at 203 Hz with an amplitude of above 4 m/s^2/sqrt(Hz). When the 50m cables were measured in the box with the conductive front panel, the amplitude drops at 203 Hz by a factor of around 3. I also compared the isolated front panel with the old setup, and found that the isolated front panel worse by a factor of just over 2 than the old setup. Therefore, I think that using the conductive front panel for the ALS Delay Line box will reduce noise and crosstalk between the cables the most. 

Attachment 1: best4.png
best4.png
Attachment 2: isolated4.png
isolated4.png
  11490   Tue Aug 11 02:40:29 2015 ericqUpdateLSC50m delay lines - Rough calibrations

Jessica will soon ELOG about some measurements suggesting that the conductive connector-ized ALS delay line enclosure is the way to go, when considering crosstalk between the delay lines. It is currently mounted and hooked up on the LSC rack, though I need to make a bunch of new SMA cables now that I think a semi-permanent arrangement has been reached. 

I did a rough re-calibration of the phase tracker output, since the increased cable delay changes the degree/Hz gain. This was done by fitting a line to a slow sawtooth FM of the SRS DS345's (1Hz rate, 10kHz deviation, 30MHz carrier). This resulted in the following calibration updates

  • ALSX: 19230 -> 13051 Hz/count, 3.4dB more sensitive

  • ALSY: 19425 -> 12585 Hz/count, 3.8db more sensitive

Again, this is a rough calibration. Nevertheless, it is not so surprising we don't get the 50m/30m = 4.4dB increase we would expect just from the lengths; the (I presume) increased cable loss matters. Also, the loss' frequency dependance is an additional reason that the phase tracker calibration is not constant over all frequencies. 

I took spectra with the arms in IR lock, but didn't see any real improvement beyond a possible dip in the floor from 100-200Hz. This doesn't surprise me too much, however, since I don't believe that we are currently dominated by electronic noises that this gain increase would help overcome. 

Last week, Koji mentioned the ALS phase noise added due to the post-cavity table motion the arm-transmitted green beams experience before hitting the beat PD. I should estimate the size of this effect for our situation. 

  11489   Tue Aug 11 02:26:46 2015 ericqUpdateASCPRC Angular FF Lives!

PRC Angular FF is back in action!

Short and sweet of it:

  • Took witness (T240 channels) and target (POP QPD) with DC coupled oplevs on. About 25 minutes of nice stationary data.
  • Downsampled everything to 32Hz, since coherence suggests subtraction only really possible from 1-5Hz. 
  • Prefiltering done by detrending and ellip(3,3,40,5Hz)
  • 4 second FIR impulse time was enough
  • Filtered target by inverse actuator TF before sending to wiener code. The only difference between this and filtering the witnesses with the actuator TF directly is an effective RMS cost function, i.e. prefiltering. 
  • Spending time tweaking IIR fitting pays off. Divided out zpk(0, [p3, p3*],1), where p3 is some well fit stack/suspension resonance, so that vectfit fits remaining portion with equal numbers of poles and zeros, guaranteeing AC coupling and 1/f rolloff to prevent noise injection
  • Quack->foton->OAF all worked fine
  • All in all, seems to work well. POPDC RMS goes down by a factor of 2 yes

  • Code used lives in /users/ericq/2015-08-PRCFF and the NoiseCancellation github repo

Fit example:


Subtraction spectra


Subtraction prediction vs. reality (positive dB is good)

Attachment 1: fitExample.png
fitExample.png
Attachment 2: FFspectra.png
FFspectra.png
Attachment 3: PITsub.png
PITsub.png
Attachment 4: YAWsub.png
YAWsub.png
  11488   Mon Aug 10 22:18:19 2015 IgnacioUpdateIOOReady to do some online mode cleaner subtraction

I'm attaching a SISO IIR Wiener filter here for reference purposes that will go online either tonight or tomorrow evening. This is a first test to convince myself that I can get this to work, MISO IIR filters are close to being ready and will soon be employed. 

This Wiener filter uses the STS-X channel as a witness and MCL as target. The bode plot for the filter is shown below,

The performance of the FIR and IIR Wiener filters and the ammount of subtraction achive for MCL is shown below,

 

Output from quack to be loaded with foton: filter.zip

K bye.

Attachment 1: stsx.png
stsx.png
Attachment 2: performance.png
performance.png
Attachment 3: filter.zip
ELOG V3.1.3-